OpenStudy (anonymous):

Identify the vertex for the graph of y = 2x2 + 8x − 3. (2, 21) (2, 17) (−2, −11) (−2, −27)

OpenStudy (anonymous):

\[y=2x ^{2}+8x-3=2\left(x ^{2}+4x+4-4 \right)-3=2\left( x ^{2}+4x+4 \right)-8-3\] \[y=2\left( x+2 \right)^{2}-11,\frac{ y+11 }{2 }=\left( x+2 \right)^{2}\] \[\left( x+2 \right)^{2}=\frac{ 1 }{2 }\left( y+11 \right)\] it is a upper parabola with vertex (-2,-11)

OpenStudy (anonymous):

@surjithayer are you still typing your response?

OpenStudy (anonymous):

Thank you so much!!!